OpenStudy (anonymous):

I need help to solve this,to represent the nth term,-7,-4,-1,2,...

OpenStudy (anonymous):

-7+3=-4 -4+3=-1 -1+3=2 .....

OpenStudy (anonymous):

hear is the whole problem Write an expression to represent the nth term of the sequence below, (explicit formula) -7,-4,-1,2,...

OpenStudy (anonymous):

\[a_{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

I know the equation but trying to solve it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

d=?

OpenStudy (anonymous):

d=3

OpenStudy (anonymous):

good

OpenStudy (anonymous):

\[a _{1}=?\]

OpenStudy (anonymous):

\[a _{1}\] is the first term of an arithmetic progression

OpenStudy (anonymous):

i solved it is a,n=3n-10

OpenStudy (anonymous):

her is another, equation Find the 47th term of each arithmetic sequence, and then write an explicit formula for each sequence. 2) 4, 6, 8, 10, 12…

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

a1=? and d=?

OpenStudy (anonymous):

a1=4 and d=2 ok?

OpenStudy (anonymous):

\[a_{n}=a _{1}+(n-1)d\] \[a_{47}=a _{1}+(47-1)d\] \[a_{47}=a _{1}+46d\]

OpenStudy (anonymous):

\[a_{47}=4+46*2\]

OpenStudy (anonymous):

the answer is a,n=50

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

46*2=?

OpenStudy (anonymous):

its a,n=-44

OpenStudy (anonymous):

no is 92

OpenStudy (anonymous):

\[a_{47}=4+92=96\]

OpenStudy (anonymous):

now the next one, 9, 5, 1, -3, -7… Find the 47th term of each arithmetic sequence, and then write an explicit formula for each sequence.

OpenStudy (anonymous):

a1=? and d=?

OpenStudy (anonymous):

a1=9 and d=(-4)

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

\[a_{47}=a _{1}+46d\]

OpenStudy (anonymous):

now substitute the value of a1 and d